### v1

@@ III Diquarks and hidden charmed tetraquarks
The masses of hidden charmed tetraquarks with strange or antistrange quark have not been calculated, and few relevant exotics have not been analyzed for the shortage of confirmed experimental data. There are some uncertainties in our calculation. First, the mixing between normal mesons and tetraquarks has not been taken into account explicitly, but was taken into account through the variation of parameters instead. The fixed parameters may vary with the components of the benchmark $X(3872)$ and $T_{cc}(3875)$, which will bring several tens MeV uncertainties to the predicted mass spectra. The boson-exchange interactions between the diquark and the antidiquark may make the predicted mass spectra lower than the predicted ones without boson-exchange interactions. So far, some charmonium-like XYZ have been suggested as the normal $c\bar{c}$ charmoniums though these assignments are not definite. In order to identify a normal charmonium or charmonium-like meson, it is time and important to take into account the mixing between the normal charmonium and the hidden-flavor charmonium-like exotics in the same time. Only when the properties of normal charmonium and hidden-flavor charmonium-like exotics are clear both in theories and in experiments, the identification could be confirmed.
@@ Acknowledgements.
This work is supported by National Natural Science Foundation of China under the grant No. 11975146. Ailin Zhang is grateful for the useful discussions with Changzheng Yuan.

### v2

@@ III Diquarks and hidden charmed tetraquarks
The masses of hidden charmed tetraquarks with a strange or antistrange quark have not been calculated, and few relevant exotics have not been analyzed because of the shortage of confirmed experimental data. There are some uncertainties in our calculation. First, the mixing between normal mesons and tetraquarks has not been taken into account explicitly, but was taken into account through the variation of parameters instead. The fixed parameters may vary with the components of the benchmark $X(3872)$ and $T_{cc}(3875)$, which will bring uncertainties of several tens of MeV to the predicted mass spectra. According to Ref. [108], boson-exchange interactions between the diquark and the antidiquark may make the predicted mass spectra lower than the predicted ones without boson-exchange interactions. So far, some charmonium-like XYZ exotics have been suggested as the normal $c\bar{c}$ charmoniums though these assignments are not definite. In order to identify a normal charmonium or charmonium-like meson, it is important to take into account the mixing between the normal charmonium and the hidden-flavor charmonium-like exotics at the same time. Only when the properties of normal charmonium and hidden-flavor charmonium-like exotics are clear both in theories and in experiments, the identification could be confirmed.
@@ Acknowledgements.
This work is supported by National Natural Science Foundation of China under Grant No. 11975146. Ailin Zhang is grateful for the useful discussions with Changzheng Yuan.